Routine check for DGS maintenance day (Feb. 4)
Check if the power supply can be controlled by BIO.
Check if the script can be started.
Check that the limit switch installed is green.
No problem (BS was in use, so it was not used this time)
Confirmed that the Release Sensor is 0. 1: Seated
ETMX, ETMY, ITMX, ITMY : No problem
Ping Test is performed.
=> Confirmation only, no recovery work was done.
PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.
Check if the temperature acquisition is working properly.
Precision air Processer
Is the temperature acquisition function working properly?